# Discrete minimizers of the interaction energy in collective behavior: a brief numerical and analytic review

### v1

@@ 4 Continuous limits
$\nu_{X_{N}}:=\frac{1}{N}\mu_{X_{N}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}^{N}}\overset{*}{\rightharpoonup}\nu\qquad\text{as $N\to+\infty$}$
@@
    One needs to show that the diameter of a minimizer $X_{N}$ stays bounded as $N\to+\infty$. As a consequence, the empirical measures $\mu_{X_{N}}/N$ have a convergent subsequence, in the weak-*** sense, to a certain probability measure $\mu$. By using the previous Gamma-convergence result (and possibly some additional uniform estimates on other properties of the minimizers) one may then deduce that $\mu$ must be a minimizer of $E$.
@@
For “strongly repulsive” potentials with $-d<b<2-d$ (but still locally integrable), regularity of minimizers can be obtained via obstacle problems (Carrillo, Delgadino, and Mellet, 2016). Explicit candidates for minimizers can be found for $-d<b<a$ with $b\leq 2$, when $a$ or $b$ are even integers (Carrillo and Huang, 2017). This is in contrast with minimizers for mildly repulsive potentials (potentials which are $\mathcal{C}^{2}$ at 00), for which minimizers must be supported on a finite number of points under quite general conditions (Carrillo, Figalli, and Patacchini, 2017). This result is true even for local minimizers in the topology of the $\infty$-Wasserstein transport distance. This applies to power-law potentials of the form (1.2) whenever $a>2$.
@@
For $b=2$ and $a\in(2,4)$, the unique minimizer (up to translations) is the uniform measure on an $n-1$-dimensional sphere of appropriate radius (Davies, Lim, and McCann, 2021). For $a\geq 4$ and $b\geq 2$ (excluding the case $(a,b)=(4,2)$) the minimizer is unique up to rigid movements, and it must be a sum of equal delta functions supported at the vertices of a regular $d$-simplex (Davies, Lim, and McCann, 2022). In this same reference, the case $(a,b)=(4,2)$ is studied. It is a special limit case in which minimizers are not unique (even up to rigid movements): minimizers may be supported on a sphere, on a finite number of points, or may be convex combinations of these possibilities. In the 1D case explicit minimizers are known for some power-law potentials of the form (1.2) (Frank, 2022).
@@
Figure 3: Numerical candidate for minimizer of the potential (1.2) with $a=2$, $b=-1$ and $10^{4}$ particles. (We thank David Wales for providing this result.) This corresponds to the “strongly repulsive” regime considered in Carrillo et al. (2016). In this regime it is conjectured that a sequence of minimizers $(X_{N})_{N\geq 2}$ must converge to a measure $\nu$ which must have a smooth density with respect to Lebesgue measure. In fact, the most likely candidate measure is given in Carrillo and Huang (2017, eq. (38)) and is of the form
$\nu(x)=C(R^{2}-|x|^{2})^{1-\frac{b+d}{2}},$
@@
with suitable constants $C,R>0$ which ensure in particular that the total mass is $1$. In the case corresponding to this figure we have $d=2$, $b=-1$, so the density of $\nu$ vanishes at the boundary. By contrast, in the case of Figure 2 the exponent $21-(b+d)/$ 2 becomes negative and the density of $\nu$ should diverge at the boundary.
@@
In addition to this, the question of stability or instability of these minimizers for several dynamical models has been investigated in a few cases; see for example Fellner and Raoul (2010a, b) (for 1D stability of sums of Dirac masses for the aggregation equation (1.4)) or Balagué et al. (2013b) (stability of spherical shells). We will also mention some results on non-isotropic potentials: see Carrillo et al. (2021) and Carrillo and Shu (2022, 2023).
@@
We will illustrate the above properties in 2D with a few numerical results for power-law potentials of the form (1.2). Let us first fix $a=2$ and look at the appearance of numerical minimizers for several values of $b$. As the exponent $b$ increases, the repulsion becomes progressively weaker. It is clearly seen in Figure 4 that minimizers become more concentrated as $b$ increases. Also, they become more concentrated on the boundary of a ball for higher $b$. For $b<0$ we are in the regime studied in Carrillo et al. (2016), so we know minimizers of $E$ should be smooth, but it is not proved that $N$-particle minimizers approach minimizers of $E$. In contrast, for $b>0$ we know from Cañizo and Patacchini (2018) that $N$-particle minimizers must approach minimizers of the continuous energy $E$, but we do not know the specific shape of these minimizers; for example, it is not proved in this range that minimizers must be regular functions on their support, or even $L^{P}$ functions for some $p\geq 1$. Reasonable candidates for minimizers of $E$ are given in Carrillo and Huang (2017), and they are all smooth functions on their support.
@@
For $a=4$ the situation is similar, but now there is a range of exponents with $b\geq 2$ where the potential $V$ is $\mathcal{C}^{2}$. This is the range studied in Davies et al. (2021) and Carrillo et al. (2017).
@@ 5 Between crystallization and continuous limits
$\frac{1}{N}\mu_{X_{N}/N^{1/s}}\overset{*}{\rightharpoonup}\mu\qquad\text{as $N\to+\infty$?}$→+∞? (5.1)
@@
There are numerical results in 1D which seem to indicate that this is the case (Carrillo et al., 2014), but this has not been systematically explored, much less the subject of any rigorous proofs.
@@ Acknowledgements
We would like to thank David Wales for several interesting discussions and help with the software GMIN, used for the main numerical calculations in this paper. We also thank J. A. Carrillo and R. Frank for some interesting discussions. The authors acknowledge support from grant PID2020-117846GB-I00, the research network RED2022-134784-T, and the María de Maeztu grant CEX2020-001105-M from the Spanish government.

### v2

@@ 4 Continuous limits
$\nu_{X_{N}}:=\frac{1}{N}\mu_{X_{N}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}^{N}}\overset{*}{\rightharpoonup}\nu\qquad\text{as $N\to+\infty$},$
@@
    One needs to show that the diameter of a minimizer $X_{N}$ stays bounded as $N\to+\infty$. As a consequence, the empirical measures $\mu_{X_{N}}/N$ converge up to a subsequence, in the weak-* sense, to a certain probability measure $\mu$. By using the previous Gamma-convergence result (and possibly some additional uniform estimates on other properties of the minimizers) one may then deduce that $\mu$ must be a minimizer of $E$.
@@
For “strongly repulsive” potentials with $-d<b<2-d$ (but still locally integrable), regularity of minimizers can be obtained via obstacle problems (Carrillo, Delgadino, and Mellet, 2016). When $a=2$ and $-d<b<2-d$, these minimizers correspond to the “fractional Barenblatt solutions” of Caffarelli and Vázquez (2011b, a); candidates for minimizers can be found for $-d<b<a$ with $b\leq 2$, when $a$ or $b$ are even integers (Carrillo and Huang, 2017). In dimension $d\geq 2$, when $a=2$ and $2-d<b<4-d$, minimizers are radially symmetric, unique, and were obtained in Carrillo and Shu (2021).
For $b=2$ and $a\in(2,4)$, the unique minimizer (up to translations) is the uniform measure on an $(n-1)$-dimensional sphere of appropriate radius (Davies, Lim, and McCann, 2021). This has recently been extended by Frank and Matzke (2023) to the case $d\geq 2$, $2\leq a\leq 4$, $b_{*}(a)\leq b\leq 2$ with $b<a$, for a certain decreasing function $b_{*}=b_{*}(a)$ which is always between $-d+3$ and $-d+4$.
@@
This is in contrast with minimizers for mildly repulsive potentials (potentials which are $\mathcal{C}^{2}$ at 00), for which minimizers must be supported on a finite number of points under quite general conditions (Carrillo, Figalli, and Patacchini, 2017). This result is true even for local minimizers in the topology of the $\infty$-Wasserstein transport distance. This applies to power-law potentials of the form (1.2) whenever $2<b<a$. For $a\geq 4$ and $b\geq 2$ (excluding the case $(a,b)=(4,2)$) the minimizer is unique up to rigid movements, and it must be a sum of equal delta functions supported at the vertices of a regular $d$-simplex (Davies, Lim, and McCann, 2022). In this same reference, the case $(a,b)=(4,2)$ is studied. It is a special limit case in which minimizers are not unique (even up to rigid movements): minimizers may be supported on a sphere, on a finite number of points, or may be convex combinations of these possibilities. In the 1D case explicit minimizers are known for some power-law potentials of the form (1.2) (Frank, 2022).
@@
Figure 3: Numerical candidate for minimizer of the potential (1.2) with $a=2$, $b=-1$ and $10^{4}$ particles. (We thank David J. Wales for providing this result.) This corresponds to the “strongly repulsive” regime considered in Carrillo et al. (2016). In this regime it is conjectured that a sequence of minimizers $(X_{N})_{N\geq 2}$ must converge to a measure $\nu$ which must have a smooth density with respect to Lebesgue measure. In fact, the most likely candidate measure is given in Carrillo and Huang (2017, eq. (38)) and is of the form
$\nu(x)=C(R^{2}-|x|^{2})_{+}^{1-\frac{b+d}{2}},$
@@
with $(\cdot)_{+}$ denoting the positive part, with suitable constants $C,R>0$ which ensure in particular that the total mass is $1$. In the case corresponding to this figure we have $d=2$, $b=-1$, so the density of $\nu$ vanishes at the boundary. By contrast, in the case of Figure 2 the exponent $21-(b+d)/$ 2 becomes negative and the density of $\nu$ should diverge at the boundary.
@@
In addition to this, the question of stability or instability of these minimizers for several dynamical models has been investigated in a few cases; see for example Fellner and Raoul (2010a, b) (for 1D stability of sums of Dirac masses for the aggregation equation (1.4)); Balagué et al. (2013b); Kolokolnikov et al. (2011) (stability of spherical shells and rings); Bertozzi et al. (2012) (behavior of patch solutions for $b=2-d$).
We will also mention some results on non-isotropic potentials: see Carrillo et al. (2021) and Carrillo and Shu (2023, 2022).
@@
We will illustrate the above properties in 2D with a few numerical results for power-law potentials of the form (1.2). Let us first fix $a=2$ and look at the appearance of numerical minimizers for several values of $b$. As the exponent $b$ increases, the repulsion becomes progressively weaker. It is clearly seen in Figures 4 and 5 that minimizers become more concentrated as $b$ increases, since repulsion becomes locally weaker. Also, they become more concentrated on the boundary of a ball for higher $b$. For $b<0$ we are in the regime studied in Carrillo et al. (2016), so we know minimizers of $E$ should be smooth, but it is not proved that $N$-particle minimizers approach minimizers of $E$. In contrast, for $b>0$ we know from Cañizo and Patacchini (2018) that $N$-particle minimizers must approach minimizers of the continuous energy $E$, but we do not know the specific shape of these minimizers; for example, it is not proved in this range that minimizers must be regular functions on their support, or even $L^{P}$ functions for some $p\geq 1$. Reasonable candidates for minimizers of $E$ are given in Carrillo and Huang (2017), and they are all smooth functions on their support.
@@
For $a=4$ the situation is similar, but now there is a range of exponents with $b\geq 2$ where the potential $V$ is $\mathcal{C}^{2}$. This is the range studied in Davies et al. (2021) and Carrillo et al. (2017). In Figure 6 we observe how the increase of $b$ causes the appearance of empty space in the middle of the particle configuration, and its further increase implies the accumulation of all particles in a ring.
@@ 5 Between crystallization and continuous limits
$\frac{1}{N}\mu_{X_{N}/N^{1/s}}\overset{*}{\rightharpoonup}\mu\qquad\text{as $N\to+\infty$?}$ (5.1)
@@
There are numerical results in 1D which seem to indicate that this is the case (Carrillo et al., 2014), but this has not been systematically explored, much less the subject of any rigorous proofs. Figure 7 shows 2D configurations of the kind $b<a<0$ and $b<-d$, $a>0$. In some of them, we appreciate certain similarities with crystallization cases (see Figure 1), but we are not aware of any works exploring their behavior. This regime may be related to scaling limits of dynamical problems with non-integrable potentials such as Oelschläger (1990), but the link is far from clear to us.
@@ Acknowledgements
We would like to thank David J. Wales for several interesting discussions and help with the software GMIN, used for the main numerical calculations in this paper. We also thank J. A. Carrillo and R. Frank for some interesting discussions. The authors acknowledge support from grant PID2020-117846GB-I00, the research network RED2022-134784-T, and the María de Maeztu grant CEX2020-001105-M from the Spanish government.